Dr. Jane is studying the growing population of a species of jellyfish in a body of water. Each year, she records the population of jellyfish. The first year, she recorded a population of 17 jellyfish. Every following year, the population increased by four times the previous year’s population. At this rate, what will be the jellyfish population in the 9th year of her study?
Answer by Boreal(15235)

The wording of this problem is somewhat ambiguous. The population increased by 4 times the previous year's population. To me, that means the first year was 17, and the 2nd year was 68+17=85. That is 17*5^1
The third year would be 85*4=425, and that is the increase, added to the start of 85. That is 17*5^2=17*25=425.
The 9th year would be 17*5^8=6,640,625
For the population to be 4 times what it was the previous year, a different picture, it would be 17*4^8=1114112